OpenStudy (bimbim2697):

a rectangle has a length of 6x^3 y^2 - 1 and a width of 3xy+2.The formaula for the perimeter of a rectangle is P=2l+2w ,where l is length and w is the width.What is the perimeter of the rectangle?Simplify your answer.

hero (hero):

Whatever the perimeter is, it will be in the form of an expression, rather than an actual number. It gives you a hint: Simplify your answer. This reveals that the answer will definitely be an expression. All you have to do is substitute l = 6x^3y^2 - 1 and w = 3xy +2 into the formula P = 2(l + w) and then simplify.

OpenStudy (bimbim2697):

i really dont get it :(

OpenStudy (bimbim2697):

thanks

hero (hero):

You'll realize that with math, the answer will not always be given to you and you will have to figure it out on your own.

hero (hero):

If someone gives you a hint, you should at least try

hero (hero):

Had to correct a mistake

hero (hero):

P = 2l + 2w P = 2(l + w) P = 2(6x^3y^2 - 1 + 3xy + 2) P = 12x^3y^2 - 2 + 6xy + 4 P = 12x^3y^2 + 6xy + 2 P = 2(6x^3y^2 + 3xy + 1) had to correct another mistake

OpenStudy (bimbim2697):

thanks

hero (hero):

yw